OpenStudy (anonymous):

Hard calculus question, using related rates: A spherical balloon is inflated with helium at the rate of 80pi ft^3/min (pi=3.14, I just don't think there a key for it). At the moment the diameter is 8ft, how fast is the diameter of the balloon changing?

OpenStudy (amistre64):

yeah, spherical ballons are almost none existent these days to compare to

OpenStudy (amistre64):

we need a formula that relates volume to diameter

OpenStudy (amistre64):

or at least to radius then we can interpolate for diameter

OpenStudy (anonymous):

my professor wrote down: v=(4pi/3)(3r^2)(r)

OpenStudy (amistre64):

there is a spurious 3 in there i think

OpenStudy (amistre64):

it should be, if i aint mistaken \[V_{sph}=\frac{4\pi}{3}r^3\]

OpenStudy (anonymous):

is that the formula?

OpenStudy (amistre64):

for the volume of a sphere; id say it is

OpenStudy (amistre64):

can you find the derviative of that for me with respect to time?

OpenStudy (amistre64):

typos not withstanding

OpenStudy (anonymous):

ok, so based on my prof's work, he put 80pi=4pi/3(3)(4^2)(r1). Sorry if this is hard to read, I'm not sure how to get the pi symbol :(

OpenStudy (amistre64):

lets forget about what your prof did; and lets see if we can discover this anew :)

OpenStudy (amistre64):

lets start by deriving the volume of a sphere; \[V=\frac{4pi}{3}r^3\] we can use this becasue it relates volume to something akin to a diameter

OpenStudy (amistre64):

can we replace r with d/2 ? since radius = half the diamter?

OpenStudy (anonymous):

oh yea, that makes sense

OpenStudy (amistre64):

\[V=\frac{4pi}{3}r^3\] \[V=\frac{4pi}{3}(d/2)^3\] \[V=\frac{4pi}{3}\frac{d^3}{8}\] \[V=\frac{4pi}{24}d^3\] \[V=\frac{pi}{6}d^3\] this gets it all talking in V and d terms

OpenStudy (amistre64):

let me know where this doesnt make sense at becasue its important that you know whats going on behind the scenes

OpenStudy (amistre64):

\[V=\frac{pi}{6}d^3\] \[\frac{d}{dt}[V=\frac{pi}{6}d^3]\] \[\frac{d}{dt}[V]=\frac{d}{dt}[\frac{pi}{6}d^3]\] \[\frac{dV}{dt}=\frac{pi}{6}\frac{d}{dt}[d^3]\] \[\frac{dV}{dt}=\frac{pi}{6}\frac{dd}{dt}\ 3d^2\] \[\frac{dV}{dt}=\frac{3pi\ d^2}{6} \frac{dd}{dt}\] \[\frac{dV}{dt}*\frac{6}{3pi\ d^2}=\frac{dd}{dt}\] \[\frac{dV}{dt}*\frac{2}{pi\ d^2}=\frac{dd}{dt}\]

OpenStudy (anonymous):

No, i think i actually understand so far, but it doesn't seem related to what I have in my notes. But your version seems much clearer.

OpenStudy (amistre64):

We now dV/dt = 80pi and the diameter is 8 ft so lets plug them in: \[\frac{dV}{dt}*\frac{2}{pi\ d^2}=\frac{dd}{dt}\] \[80pi*\frac{2}{pi\ 8^2}=\frac{dd}{dt}\] \[80*\frac{2}{8^2}=\frac{dd}{dt}\] \[\frac{80*2}{8*8}=\frac{dd}{dt}\] \[\frac{10*2}{8}=\frac{dd}{dt}\] \[\frac{10}{4}=\frac{dd}{dt}=\frac{5}{2}\]

OpenStudy (amistre64):

there are more than one way to resolve this problem; i perfer to do it the easiest way :)

OpenStudy (anonymous):

Wow, thats amazing. Are you a genius or something, lol?

OpenStudy (amistre64):

my kids would say: or something ;)

OpenStudy (anonymous):

Do you teach math?

OpenStudy (amistre64):

depends, did I teach you?

OpenStudy (anonymous):

Just now, for sure :)

OpenStudy (amistre64):

then maybe :) good luck with it all

OpenStudy (anonymous):

thanks, I've been so stressed since i have an exam tmrw. but thanks for taking the time to help me out :)